RC meeting and Run Plan, Saturday December 18th

From SBS wiki
Jump to: navigation, search


RC Meeting Information

Run Coordinator: Bogdan Wojtsekhowski

Location: Counting House Conference Room

Time: was canceled due to weekend

Remote connection link: https://bluejeans.com/256205934

Toll-Free Number (U.S.& Canada): 1-888-240-2560

Bluejeans Meeting ID: 256 205 934

Upcoming Shifts Which Need Filled

Shifts in January need help - 195 openings to fill

Today's estimate: Optic beam Tue. Jan 11, "Bonus beam" Jan 12-16. Official beam Jan 17

Accelerator and Hall Status

  • Preparation for beam off period
  • End of beam: Tue. Dec. 21 at 7:00 AM
  • Tuesday 12/21 at 8pm expected test of the Access Control System for one hour

To be Done During Extended Down Time

  • Replace J0 INFN plane with UVa plane UV#4 and related electronics
  • Spectrometers positions for SBS-14 kinematics
  • Cosmic tests of HCal and BBcal
  • Study of GRINCH

Run Plan

  • Low current runs on Dummy for GEM efficiency test
  • Continue with "Cycle 23"
  • Continue the cycle pattern

Opportunistic Access Completed

  • None

Resolved Issues

  • Oscilloscope on SL monitor for instant check of the delays
  • GEM sector loss => reduce HV by 25 V on all GEMs

Unresolved Issues

  • NEW: HCAL ROC17 ERROR for last 12+ hours
  • Occasional timing shifts of ~70 ns
  • GEM efficiency analysis.
    • Use of the Y-tgt cut for determination of the actual efficient of the detector.

Subsystem Updates

  1. DAQ general
    1. Alex update
  2. HCal
    1. F1 errors (HCal sync errors)... Don't ignore!! End, Reset(twice?), Configure, Download. If that doesn't work, call DAQ on call.
    2. Prints of good HCal plots for LH2 and LD2 target Sebatian put on the SL table.
  3. BBCal
    1. no update. Problem with BBcal HV for last few hours.
  4. Timing Hodoscope
    1. no update
  5. GRINCH
    1. Maria presented the ADC spectra with a single PE peak from actual GMn data!
  6. GEMS
    1. DAQ update
    2. HV and efficiency status: Kondo will take data with dummy target at two HV values.
  7. Beamline
    1. David continue data analysis.
  8. Replay analysis
    1. Andrew update and results
    2. Massive data replay: script preparation is in progress.